The human body can survive a negative sudden stop if the magnitude of the acceleration is less than 250m/s2. If you are in an automobile accident with an initial speed of 105km/h and are stopped by an airbag that inflates from the dashboard, over what distance must the airbag stop you for you to survive the crash?



1
Expert's answer
2021-03-29T09:04:01-0400
"v^2=v_0^2+2ad,""0=v_0^2+2ad,""d=\dfrac{-v_0^2}{2a},""d=\dfrac{-(29.16\ \dfrac{m}{s})^2}{2\cdot(-250\ \dfrac{m}{s^2})}=1.7\ m."